OpenStudy (anonymous):

write an equation for the line with x-intercept 3 and y-intercept 2/3 y=mx+b 2y/3 =1/3x+b but how do i find my b? show all working please

OpenStudy (anonymous):

y=-2x/9+2/3

OpenStudy (saifoo.khan):

whn x intercepts y = 0 so 2y/3 =1/3x+b 2(0)/3 =1/3(3)+b 0 = 1 +b b = -1

OpenStudy (saifoo.khan):

now whts that?

OpenStudy (anonymous):

m=-b/x

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

no thats not how we solve it my apologies

OpenStudy (anonymous):

this is the double intercept so we use the formula x/a+y/b=1 where a is the x-intercept and b is the y-intercept

OpenStudy (anonymous):

x/3+3y/2=1 y=-2x/9+2/3

OpenStudy (anonymous):

thanks still saifoo.khan

OpenStudy (radar):

You are provided with 2 points P1(0,2/3) and P2(2,0) Use the formula for slope m=(y2-Y1)/(x2-x1)=(0-2/3)/(2-0)=(-2/3)/2=-1/3

OpenStudy (radar):

Slope is as shown -1/3 Now pick one of the points and use the formula y=mx+b using the point P1(0,-2/3) plug in those values -2/3=-1/3(0)+b b=-2/3 y=-2/3x-2/3